OpenStudy (anonymous):

Two six sided dice (one is red, one is green) are rolled. Determined the following probabilities: 1) What is the probability that the red die is even and the green die is odd? That is, P(R=even∩G=odd) 2)What is the probability that the sum of the red and green values is 8 . 3) What is the probability that the sum of the red and green values is <7 ? That is, P(R+G<7). 4) What is the probability that red value is ≤ green value? That is, P(R≤G) 5) What is the probability that red value ≤ green and the green value is <6 ? That is, P(R≤G∩G<6)

OpenStudy (anonymous):

1 is .25 do you know why?

OpenStudy (anonymous):

2 is 1/36

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no sorry 7/36

OpenStudy (anonymous):

ok for number 1 do you know that if you have two independent probabilities you multiply them to find their combined probability?

OpenStudy (anonymous):

so the P(even dice roll)=.5 and P(odd dice roll)=.5 so .5•.5=.25

OpenStudy (anonymous):

there like 2 dice(6*6=36)

OpenStudy (anonymous):

i see, so is 1/4

OpenStudy (anonymous):

\[\frac{ 3 }{ 6 } \times \frac{ 3 }{ 6 }=?\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

that solve the first part, is 1/4

OpenStudy (anonymous):

yea but i would use decimals unless it asks for fractions

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ok so #2

OpenStudy (anonymous):

there are 36 total combinations and only 7 add up to 8

OpenStudy (anonymous):

so 7/36 would be the probability

OpenStudy (anonymous):

k

OpenStudy (anonymous):

sorry wait i didnt account for 1 so its 8/36 which can be simplified

OpenStudy (anonymous):

1,7 2,6 3,5 4,4 4,4 5,3 6,2 1,7

OpenStudy (anonymous):

those are the combos that add up to 8 for the red and green

OpenStudy (anonymous):

16/36?

OpenStudy (anonymous):

well no because there are 16 numbers but they come in pairs

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

ya u are right

OpenStudy (anonymous):

{1,7} , so on

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok so #2 is?

OpenStudy (anonymous):

2. 26,35,44,53,62 P=5/36

OpenStudy (anonymous):

^^^^^

OpenStudy (anonymous):

but its 6 not 5 because th 4,4 repeats

OpenStudy (anonymous):

that solve second question

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

correct.6/36

OpenStudy (anonymous):

so #3 do you know what to do??

OpenStudy (anonymous):

finding a numbers that is less that 7

OpenStudy (anonymous):

yea but combos like {1,5} so on

OpenStudy (anonymous):

but not =7 its <7

OpenStudy (anonymous):

ya i notice that

OpenStudy (anonymous):

12/36

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

11,12,13,14,15,21,22,23,24,31,32,33,41,42,51first is G,second is Red same is first Red,second is Green P=30/36=5/6

OpenStudy (anonymous):

try to make a diagram or draw it

OpenStudy (anonymous):

already did

OpenStudy (anonymous):

i miss some numbers

OpenStudy (anonymous):